16:00 Beam down due to injector work.
16:15 Alexandre and Jack found that Q2 has it's set current at zero No alarm sounded in the counting house.
16:20 Called MCC to get an update on the beam status. They say 50 minutes to beam return.
16:40 Jack took a look at why HRS-L Q2 was off and no one noticed. It turns out that if the power supply is switched off the current ramps down before switching the supply off AND the alarm handler no longer catches the error. So we didn't have an alarm in the counting house as a result of the current being out of tolerance.
17:12 BigBite E and dE HV tripped off. Reset from GUI.
17:25 Alexandre requested a controlled access to make a change to the DAQ. MCC was called. Tried to page Heidi but the page wouldn't go through.
17:34 It seems beam has been restored to Hall C. We are waiting on our escort for the hall access.
17:51 Requested Beam Permit.
18:08 Back in beam permit, but no beam delivery yet.
18:13 MCC reports that they are about to deliver beam to the hall.
16:18 Beam looks good. We'll restart production at 5uA with run number 2923, Run 2922 was spot and scaler check.
18:23 BB HV trip. Reset and restarted run.
19:00 So far so good. Since the return of the beam about 40 minutes ago, things have been pretty quiet. The back room is looking at their DAQ changes from earlier in the shift, but have not yet reported. The detectors look OK, although, the Cerenkov looks a little strange. The HV is on. We've consulted with Alexandre (who happens to be in the back room) and we're checking the T5's to see if that's the issue.
21:11 Requested controlled access to add cables to detectors; which already had cables, but need more. I believe we are working on delay, also no pun intended.
23:00 Still down in the hall.
23:58 The shift ends quietly with a crew down in the hall fixing the delay problems. We had no further difficulties from either of the Q2 magnets and only occasional trips on BigBite, all of which were successfully reset. The beam has been stable since recovery at mid-shift. The beam positions were good immediately upon recovery and remained stable for the duration of the data-taking time during this shift. The detectors looked OK for replayed data (except, of course for the timing problems). Run numbers for this shift are listed in a separate entry.